• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Seonho Choi

    Log entry time 16:00:14 on December 9,2007

    Entry number 219406

    This entry is a followup to: 219402

    keyword=Update on He level crisis in L-dipole

    Although that He seems to flow in, the level is not changing much. Sometimes going under 41% again. Cryo coordinator has been contacted and he will come and see if there is any valve problem or other in the Hall. It will be in about half an hour.


    FIGURE 1